orders heparin 15,000 U sQ. Heparin is available in a dosage of 20,000 e the amount that must be administered. Antibiotics are available in powder form must be form for injection; however, antibiotics in in water or another diluent. Antibiotic powder should be diluted so the desired dose is solution. 21 mL or mL of A physician orders 400,000 U of penicillin V PO 16h for a staph infection. The dosage on hand contains 800,000/5 mL. Calculate the number of milliliters to administer the doctors order. 22. Intravenous Solutions 23. True or False? Replacement fluids help maintain normal electrolyte and fluid balances. 24. A 500 mL quantity of DIOW contains how many grams of dextrose? 25. A physician prescribes D5W 1 L with heparin 40,000 U IV at 4.200 U/hr. Calculate Flow and Drop Rates 26. Rantidine is infused in a 0.5 L bag 1/2 NS for ulcer treatment. It runs at 30 drops the flow rate in ml per hour. per min with the infusion device set to deliver 16 drops per ml. How long will the infusion take to complete? (Round to the nearest minute.) A physician orders an IV of cefazolin (Ancef) 1g/50 mL IVPB g6h. The infusion can be given in 20 minutes. What rate will you set on the IV pump (mL/min)? 27. An IV infusion order is written for 0.3 L of D5W/0.45% NS to run over 12 hours. The drop factor on the IV is set to 12 gtt/mL. What is the flow rate in drops per minute? Hint: Make sure your equation includes conversion factors for volume and time so that the answer is in gtt/min. 28. Calculate the flow rate for an IV of 0.6 L of Lactated Ringers solution in 24 hours if the drop factor for the tubing is 12 gtts/mL. Hint: Make sure your equation includes conversion factors for volume and time so that the answer is in gtt/min. 29. 12
Answer #1

Q20: answer: Liquid, Dilute

Q21: 100ml

Q22: Physician Order =400000U

Available= 800000U/ 5ml

So addictions half dose of above= 400000U/ 2.5ml

Q23: True

Q24: 10g dextrose

Q25:

Formula= amount of fluid× drop factor/ hour× minutes

= 1000× 60/ 1× 60

= 1000drops / minutes

1ml quality to 20 drops

= 1000/ 20

= 50ml per hour.

Q26:

500ml infusion runs by 30drops per minutes.

500ml infusion runs by 16drops per minutes

1ml equal to 20 drops

0.5ml = 10drops

0.25ml= 5drops

0.30ml= 6drops

0.8ml= 16drops

Per minute 0.8ml

60minutes = 60× 0.8ml= 48ml

10 hours= 48× 10ml= 480ml

30minutes= 20ml

So around 10hour and 30minutes to complete 500ml.

Q27:

50× 15/ 0.33× 60

4.125 drops / minutes

Q28:

Formula= amount of fluid× dropped factor/ hour× minutes

= 300× 12/ 12× 60

= 5 drops / minutes

Q29:

Formula= amount of fluid× drop factor/ time× minute

= 600× 12/24× 60

5 drops per minute